# Let It Flow: Simultaneous Optimization of 3D Flow and Object Clustering

### v1

@@ Abstract
We study the problem of self-supervised 3D scene flow estimation from real large-scale raw point cloud sequences, which is crucial to various tasks like trajectory prediction or instance segmentation. In the absence of ground truth scene flow labels, contemporary approaches concentrate on deducing optimizing flow across sequential pairs of point clouds by incorporating structure based regularization on flow and object rigidity. The rigid objects are estimated by a variety of 3D spatial clustering methods. While state-of-the-art methods successfully capture overall scene motion using the Neural Prior structure, they encounter challenges in discerning multi-object motions. We identified the structural constraints and the use of large and strict rigid clusters as the main pitfall of the current approaches and we propose a novel clustering approach that allows for combination of overlapping soft clusters as well as non-overlapping rigid clusters representation. Flow is then jointly estimated with progressively growing non-overlapping rigid clusters together with fixed size overlapping soft clusters. We evaluate our method on multiple datasets with LiDAR point clouds, demonstrating the superior performance over the self-supervised baselines reaching new state of the art results. Our method especially excels in resolving flow in complicated dynamic scenes with multiple independently moving objects close to each other which includes pedestrians, cyclists and other vulnerable road users. Our codes will be publicly available.
Figure 1: Comparison of the proposed method with self-supervised competitors on Argoverse2 Dataset [3]. Our method is able to distinguish the different motion patterns and separate objects, while other methods tend to under-segment objects and fit incorrect rigid motion.
1 Introduction
@@
Self-supervision in 3D flow estimation always stems from strong prior assumptions about the scene structure. Many existing approaches, as exemplified by works [1, 11, 14, 19, 33], assume that the underlying scene consists of a static background and several rigid spatially compact objects, the motion of which satisfies kinematic and dynamic constraints. State-of-the-art methods [11, 19, 33, 7, 36] typically construct spatially compact clusters [9], that are assumed to correspond to rigidly moving objects as a first step and then apply rigid flow regularization for these clusters. Such premature clustering inevitably suffers from under-segmentation (i.e. merging several rigid objects into a single cluster) and over-segmentation (i.e. disintegration of a single rigid object to multiple clusters). To partially suppress the under-segmentation issue, the outlier rejection technique has recently been proposed [36]. Nevertheless, the quality of the initial clustering remains the main bottleneck, which prevents estimating the accurate flow, especially in situations where multiple small objects are moving independently close to each other. In contrast to existing approaches, we generate many small overlapping spatio-temporal rigid-cluster hypotheses and then jointly optimize the flow with the rigid-body segmentation. We quantitatively and qualitatively demonstrate that such an approach mitigates the over and under-segmentation issues and consequently yields superior results, especially on dynamic objects; see Figure 1 for the comparison.
@@
Figure 2: Outline of the proposed losses: The left image shows pointcloud represented by black crosses that contains two vertical rigid objects. There is no spatial clustering that would segment them correctly; therefore any independent flow estimation will be strongly biased by the incorrect clustering. In contrast, we cover the pointcloud by (i) non-overlapping hard rigid clusters and (ii) overlapping soft rigid clusters. The right image demonstrates losses, visualized by springs, that are used for the flow estimation. The resulting flow is used for merging the hard clusters on spatio-temporal domain. The procedure is repeated until convergence. The resulting hard rigid clusters delivers rigid object segmentation.
@@
In the proposed self-supervised framework, we employ two losses that are visualized using the analogy of a mechanical machine [^1] Since both losses minimize the L2-norm of some quantities, we can visualize their influence on the estimated flow by ideal springs. In the ideal spring, the total conserved energy is proportional to the square of its deformation; therefore, the value of L2-loss corresponds to its energy. with springs, see Figure 2. The equilibrium of this machine corresponds to the globally optimal flow. The machine consists of two consecutive pointclouds in time (black and red crosses) which are not allowed to move, flow arrows attached to blue points by swivel telescopic joints, and two losses (rigid and distance) represented by springs. The scene depicts two rigid vertical objects that are moving in different directions. We employ standard distance loss (red springs), which attracts the flow of black points toward the corresponding red points from the consecutive pointcloud. Since objects in black pointcloud are close to each other, there is no spatial clustering that would separate them correctly. In contrast to existing methods, we instead introduce overlapping clusters (blue/green/magenta regions), within which the rigidity is encouraged through rigidity loss (blue/green/magenta springs). This rigid loss constructs springs among the arrow end-points in each source rigid-clusters, which prevents the flow from deforming the rigid cluster. The color of the springs matches the color of the cluster.
@@
We also enable outlier rejection through spectral clustering as proposed in [36]. In terms of the mechanical machine analogy, the intuition of outlier rejection is that it adjusts the stiffness of springs in order to weaken the springs that connect outlier points (i.e. the points, the motion of which is non-rigid with respect to the rest of the cluster). To further regularize the flow, a structural regularization through Neural prior [20, 36, 21, 7] has also been proposed. We identified that the Neural prior regularizer helps mostly with static scenes flow estimation with a small number of dynamic objects; however, it over-regularizes the flow in highly dynamic scenes due to the limited capacity of the Neural-prior network. Therefore, we decided to drop this regularization in our proposed approach.
@@
2 Related Work
Supervised scene flow.
@@ 2 Related Work
In the early stages of learning-based 3D scene flow estimation, methodologies relied on synthetic datasets [12, 26, 38, 25] for preliminary training. FlowNet3D [24], assimilated principles from FlowNet [13] and underwent comprehensive supervised training, employing L2 loss in conjunction with ground-truth flow annotations. BiPFN [6] orchestrated the bidirectional propagation of features from each point cloud, thereby enriching the representation of individual points. FLOT [31] introduced an innovative correspondence-based network, transforming the initial flow estimation to optimal transport estimation problem, succeeded by meticulous flow refinement through trained convolutions. The recently unveiled IHNet architecture [39] proposes an iterative hierarchical network, steered by high-resolution inferred information and estimating flow in a coarse-to-fine manner.
Self-supervised scene flow.
@@
Recent methodologies have sidestepped the requirement for ground-truth flow by embracing self-supervised learning. Among the early adopters of this approach, PointPWC-Net [41] presented a wholly self-supervised methodology, combining nearest-neighbors and Laplacian losses. Similarly, in [28], a self-supervised nearest neighbor loss was implemented, ensuring the cycle consistency between forward and reverse scene flow. Najibi et al. [29] employ self-learned flow within an automated pseudo-labeling pipeline, designed for training an open-set object detector and trajectory predictor. RigidFlow [19] formulates a methodology for generating pseudo scene flow within the domain of self-supervised learning. This approach relies on piecewise rigid motion estimation applied to collections of pseudo-rigid regions, identified through the supervoxels method [23]. Flowstep3D [16] undertakes flow estimation through local and global feature matching utilizing correlation matrices and iterative warping. The method known as SCOOP [17] adopts a hybrid framework of correspondence-based feature matching, coupled with a flow refinement layer employing self-supervised losses. Shen et al. [32] seamlessly integrated superpoint generation directly into their model for iterative flow refinement. Compared to supervised and self-supervised approaches, our method do not require any labeled or unlabeled training data as it is optimization-based.
Direct flow optimization.
@@
The Graph Prior approach, as introduced by [30], is rooted in pure optimization, enabling flow estimation without the need for training data. The Neural Prior[20] illustrates that flows can be optimized by incorporating a structure-based prior within the network architecture. The central objective in [21] revolves around expediting the Neural Scene Flow [21]. Li et al. identify the Chamfer distance as a computational bottleneck and use the Distance Transform as a surrogate correspondence-free loss function. The methodology outlined in Scene Flow via Distillation by [35] embodies a direct distillation framework. This innovative approach utilizes a label-free optimization method to generate pseudo-labels, which are subsequently employed to supervise a feed-forward model to achieve better speed/performance ratio. The two newest optimization-based methods, the MBNSFP [36] and Chodosh [7] adopt Neural Prior architecture and enhanced it with spatial consistency regularization [36] and post-process cluster rigidity [7], respectively. We also use rigidity regularization, but we do not limit our method to fixed initial clusters as [36, 7]. Instead, we jointly optimize initial cluster and flow to progressively merge and enlarge clusters while enforcing rigidity on them. Another difference to aforementioned methods, is the absence of Neural Prior as structural regularization, which plays crucial role for populated dynamic scenes.
Rigidity Regularization.
@@
Given the presence of multiple local minima in the conventional self-supervised Chamfer distance loss, the introduction of regularization mechanisms becomes imperative to attain physically plausible flows—ones that faithfully trace the motion of rigid structures and objects. Implicit rigidity regularization is enforced through a robust model prior in [20, 2, 42]. The introduction of weak supervision, in the form of ego-motion and foreground segmentation, has demonstrated its capacity to provide object-level abstraction for estimating rigid flows [11]. FlowStep3D [16] imposed constraints on the source point cloud to preserve its Laplacian when warped by the predicted flow, constituting the smoothness loss. SLIM [1] incorporated the smoothness loss into the process of learning network weights and SCOOP [17] integrated the smoothness loss within the optimization-based flow refinement layer. In [5], a novel global similarity measure takes the form of a second-order spatial compatibility measure on consensus seeds, that serve as input to a weighted SVD algorithm, ultimately producing global rigid transformation. Contrary to [5], we use spatial consistency mechanism for rigidity regularization, but for multiple objects.
3 Method
@@
The main goal is the self-supervised estimation of the flow field $F$ between two consecutive point clouds $P$ and $Q$. Point clouds consist of 3D points $\mathbf{p}_{i}\in P\subset\mathbb{R}^{3}$ and $\mathbf{q}_{i}\in Q\subset\mathbb{R}^{3}$. Since real-world scenes consist mostly of a static background, we follow a good practice of compensating the ego-motion [7] first before running our method. We calculate the ICP [37] as an estimate of ego-motion and use it to transform the source point cloud $P$. In the rest of this section, we focus on the estimation of the remaining flow, which corresponds to the motion between transformed source pointcloud $P$ and target pointcloud $Q$. Consequently, the flow of point $\mathbf{p}_{i}\in P$ is a 3D vector $\mathbf{f}_{i}\in F\subset\mathbb{R}^{3}$ that is non-zero only on dynamic objects if the ego-motion is correctly estimated.
3.1 Method overview
@@
The method is summarized by pseudo-code in Algorithm 1. The proposed method first initialize two sets of clusters: hard and soft. Hard rigid clusters are non-overlapping small clusters that are assumed to cover only a single rigid object or its part. Each rigid cluster $H\in\mathcal{H}$ is a small compact cluster delivered through spatial-temporal segmentation on $P$ including the adjacent temporal point clouds. In contrast, the soft rigid clusters are small overlapping clusters that are expected to overflow into neighboring objects. Each point $\mathbf{p}\in P$ is associated with the one soft cluster $S\in\mathcal{S}$, which is defined as the set of its $k$ nearest neighbours $N_{k}(\mathbf{p},P)$ from the pointcloud $P$. Given these two sets of clusters, we optimize the flow to minimize (i) hard rigidity loss on hard clusters, (ii) soft rigidity loss on soft clusters and (iii) distance loss on all points. Both rigidity losses enforce rigid flow on points within the cluster. The main difference is that the soft rigidity loss allows for outlier rejection. All losses are detailed in the following paragraphs. Once the optimized flow is available, we merge rigid clusters in $P$, whose flow goes into the same rigid cluster in $Q$. Since the algorithm is iteratively called on all consecutive pairs of pointclouds, the rigidity is propagated throughout the temporal domain.
@@
    Optimize
    $\mathcal{L}(F)=\alpha\sum_{\mathbf{f}\in F}\mathcal{L}_{\textrm{dist}}(\mathbf{f})+\beta\sum_{H\in\mathcal{H}}\sum_{\mathbf{f}\in H}\mathcal{L}_{\textrm{hard}}(\mathbf{f})+\gamma\sum_{S\in\mathcal{S}}\sum_{\mathbf{f}\in S}\mathcal{L}_{\textrm{soft}}(\mathbf{f}),$
    where $\alpha,\beta$ and $\gamma$ are hyper-parameters of the proposed method.
@@
    Repeat from 3 until convergence or reaching maximum number of iterations.
3.2 Distance loss
@@
$\mathcal{L}_{\mathrm{dist}}(\mathbf{f})=\sum_{i\in P}\|\mathbf{p}_{i}+\mathbf{f}_{i}-N_{1}(\mathbf{p}_{i}+\mathbf{f}_{i},Q)\|$ (1)
@@
3.3 Hard rigidity loss

### v2

@@ Abstract
We study the problem of self-supervised 3D scene flow estimation from real large-scale raw point cloud sequences. The problem is crucial to various automotive tasks like trajectory prediction, object detection, and scene reconstruction. In the absence of ground truth scene flow labels, contemporary approaches concentrate on deducing and optimizing flow across sequential pairs of point clouds by incorporating structure-based regularization on flow and object rigidity. The rigid objects are estimated by a variety of 3D spatial clustering methods. While state-of-the-art methods successfully capture overall scene motion using the Neural Prior structure, they encounter challenges in discerning multi-object motions. We identified the structural constraints and the use of large and strict rigid clusters as the main pitfall of the current approaches, and we propose a novel clustering approach that allows for a combination of overlapping soft clusters and non-overlapping rigid clusters. Flow is then jointly estimated with progressively growing non-overlapping rigid clusters together with fixed-size overlapping soft clusters. We evaluate our method on multiple datasets with LiDAR point clouds, demonstrating superior performance over the self-supervised baselines and reaching new state-of-the-art results. Our method excels in resolving flow in complicated dynamic scenes with multiple independently moving objects close to each other, including pedestrians, cyclists, and other vulnerable road users. Our codes are publicly available on https://github.com/ctu-vras/let-it-flow.
Index Terms:
Autonomous Driving, LiDAR, 3D Scene Flow, Object Clustering
Figure 1: Performance comparison of the proposed method with self-supervised competitors on Argoverse2 Dataset [3]. Our method is able to distinguish the different motion patterns and separate objects, while other methods tend to under-segment objects and fit incorrect rigid motion. The qualitative example comes from Waymo [15] dataset.
I Introduction
@@
Self-supervision in 3D flow estimation always stems from strong prior assumptions about the scene structure. Many existing approaches, as exemplified by works [1, 11, 14, 19, 34], assume that the underlying scene consists of a static background and several rigid spatially compact objects, the motion of which satisfies kinematic and dynamic constraints. State-of-the-art methods [11, 19, 34, 7, 37] typically construct spatially compact clusters [9], that are assumed to correspond to rigidly moving objects as a first step, and then they apply rigid flow regularization to these clusters. Such premature clustering inevitably suffers from under-segmentation (i.e., merging several rigid objects into a single cluster) and over-segmentation (i.e., disintegration of a single rigid object into multiple clusters). To partially suppress the under-segmentation issue, the outlier rejection technique has recently been proposed [37]. Nevertheless, the quality of the initial clustering remains the main bottleneck, which prevents estimating the accurate flow, especially in situations where multiple small objects move independently close to each other. In contrast to existing approaches, we generate many small overlapping spatio-temporal rigid-cluster hypotheses and then jointly optimize the flow with the rigid-body segmentation. We quantitatively and qualitatively demonstrate that such an approach mitigates the over and under-segmentation issues and consequently yields superior results, especially on dynamic objects; see Figure 1 for the comparison.
@@
Figure 2: Outline of the proposed losses: The left image shows a point cloud, represented by black crosses, containing two vertical objects; the silhouettes were added for visualization. No spatial clustering would segment the objects correctly; therefore, any independent flow estimation will be strongly biased by the incorrect clustering. In contrast, we cover the point cloud by (i) non-overlapping hard rigid clusters and (ii) overlapping soft rigid clusters. The right image demonstrates losses, visualized by springs, used for the flow estimation. The resulting flow is used to merge the hard clusters in the spatio-temporal domain. The procedure is repeated until convergence. The resulting hard rigid clusters deliver rigid object segmentation.
@@
In the proposed self-supervised framework, we employ two losses that are visualized using the analogy of a mechanical machine [^1] Since both losses minimize the L2-norm of some quantities, we can visualize their influence on the estimated flow by ideal springs. In the ideal spring, the total conserved energy is proportional to the square of its deformation; therefore, the value of L2-loss corresponds to its energy. with springs, see Figure 2. The equilibrium of this machine corresponds to the globally optimal flow. The machine consists of two consecutive point clouds in time (black and red crosses), which are not allowed to move, flow arrows attached to blue points by swivel telescopic joints, and two losses (rigid and distance) represented by springs. The scene depicts two rigid vertical objects that are moving in different directions. We employ standard distance loss (red springs), which attracts the flow of black points toward the corresponding red points from the consecutive point cloud. Since objects in the black point cloud are close to each other, no spatial clustering would separate them correctly. In contrast to existing methods, we instead introduce overlapping clusters (blue/green/magenta regions), within which the rigidity is encouraged through rigidity loss (blue/green/magenta springs). This rigid loss constructs springs among the arrow end-points in each source rigid-clusters, which prevents the flow from deforming the rigid cluster. The color of the springs matches the color of the cluster.
@@
We also enable outlier rejection through spectral clustering as proposed in [37]. In terms of the mechanical machine analogy, the intuition of outlier rejection is that it adjusts the stiffness of springs in order to weaken the springs that connect outlier points (i.e., the points, the motion of which is non-rigid with respect to the rest of the cluster). To further regularize the flow, a structural regularization through Neural prior [20, 37, 21, 7] has also been proposed. We identified that the Neural prior regularizer helps mainly with scene flow estimation in static scenes with a small number of dynamic objects; however, it over-regularizes the flow in highly dynamic scenes due to the limited capacity of the Neural-prior network. Therefore, we decided to drop this regularization in our proposed approach.
@@
II Related Work
Supervised scene flow
@@ II Related Work
In the early stages of learning-based 3D scene flow estimation, methodologies relied on synthetic datasets [12, 27, 39, 26] for preliminary training. FlowNet3D [25], assimilated principles from FlowNet [13] and underwent comprehensive supervised training, employing L2 loss in conjunction with ground-truth flow annotations. BiPFN [6] orchestrated the bidirectional propagation of features from each point cloud, thereby enriching the representation of individual points. FLOT [32] introduced an innovative correspondence-based network, transforming the initial flow estimation to optimal transport estimation problem, succeeded by meticulous flow refinement through trained convolutions. The recently unveiled IHNet architecture [40] proposes an iterative hierarchical network steered by high-resolution inferred information and estimating flow in a coarse-to-fine manner.
Self-supervised scene flow
@@
Recent methodologies have sidestepped the requirement for ground-truth flow by embracing self-supervised learning. Among the early adopters of this approach, PointPWC-Net [42] presented a wholly self-supervised methodology, combining nearest-neighbors and Laplacian losses. Similarly, in [29], a self-supervised nearest neighbor loss was implemented, ensuring the cycle consistency between forward and reverse scene flow. Najibi et al. [30] employ self-learned flow within an automated pseudo-labeling pipeline, designed for training an open-set object detector and trajectory predictor. RigidFlow [19] formulates a methodology for generating pseudo scene flow within the domain of self-supervised learning. This approach relies on piecewise rigid motion estimation applied to collections of pseudo-rigid regions, identified through the supervoxels method [24]. Flowstep3D [16] undertakes flow estimation through local and global feature matching utilizing correlation matrices and iterative warping. The method known as SCOOP [17] adopts a hybrid framework of correspondence-based feature matching coupled with a flow refinement layer employing self-supervised losses. Shen et al. [33] seamlessly integrated superpoint generation directly into their model for iterative flow refinement. Compared to supervised and self-supervised approaches, our method does not require any labeled or unlabeled training data as it is optimization-based.
Direct flow optimization
@@
The Graph Prior approach, as introduced by [31], is rooted in pure optimization, enabling flow estimation without the need for training data. The Neural Prior[20] illustrates that flows can be optimized by incorporating a structure-based prior within the network architecture. The central objective in [21] revolves around expediting the Neural Scene Flow [21]. Li et al. identify the Chamfer distance as a computational bottleneck and use the Distance Transform as a surrogate correspondence-free loss function. The methodology outlined in Scene Flow via Distillation by [36] embodies a direct distillation framework. This innovative approach utilizes a label-free optimization method to generate pseudo-labels, which are subsequently employed to supervise a feed-forward model to achieve a better speed/performance ratio. The two newest optimization-based methods, the MBNSFP [37] and Chodosh [7] adopt Neural Prior architecture and enhanced it with spatial consistency regularization [37] and post-process cluster rigidity [7], respectively. We also use rigidity regularization, but we do not limit our method to fixed initial clusters as [37, 7]. Instead, we jointly optimize the initial cluster and flow to merge and enlarge clusters while enforcing rigidity on them progressively. Another difference between our method and the aforementioned methods is the absence of Neural Prior as structural regularization, which allows us to capture correct flows of multiple mutually-exclusive object motions.
Rigidity Regularization
@@
Given the presence of multiple local minima in the conventional self-supervised Chamfer distance loss, introducing regularization mechanisms becomes imperative to attain physically plausible flows—ones that faithfully trace the motion of rigid structures and objects. Implicit rigidity regularization is enforced through a robust model prior in [20, 2, 43]. The introduction of weak supervision, in the form of ego-motion and foreground segmentation, has demonstrated its capacity to provide an object-level abstraction for estimating rigid flows [11]. FlowStep3D [16] imposed constraints on the source point cloud to preserve its Laplacian when warped by the predicted flow, constituting the smoothness loss. SLIM [1] incorporated the smoothness loss into the process of learning network weights and SCOOP [17] integrated the smoothness loss within the optimization-based flow refinement layer. In [5], a novel global similarity measure takes the form of a second-order spatial compatibility measure on consensus seeds that serve as input to a weighted SVD algorithm, ultimately producing global rigid transformation. Contrary to [5], we use a spatial consistency mechanism for rigidity regularization but for multiple objects. ICP-flow [23] performs hierarchical clustering with histogram initialization for ICP registration per object to mitigate registration local minima. Such clustering usually suffers from the under-segmentation of objects. We differ from [23] by utilizing per-point soft registration and specifically addressing the under-segmentation by including the clustering in the optimization process. We also differ in performing the clustering in spatial-temporal domain, rather than geometrical space, and we do not require matching the clusters in two consecutive frames.
III Method
@@
The main goal is the self-supervised estimation of the flow field $F$ between two consecutive point clouds $P$ and $Q$. Point clouds consist of 3D points $\mathbf{p}_{i}\in P\subset\mathbb{R}^{3}$ and $\mathbf{q}_{i}\in Q\subset\mathbb{R}^{3}$. Since real-world scenes consist mostly of a static background, we follow a good practice of compensating the ego-motion [7] first before running our method. We calculate the ICP [38] as an estimate of ego motion and use it to transform the source point cloud $P$. In the rest of this section, we focus on the estimation of the remaining flow, which corresponds to the motion between transformed source point cloud $P$ and target point cloud $Q$. Consequently, the flow of point $\mathbf{p}_{i}\in P$ is a 3D vector $\mathbf{f}_{i}\in F\subset\mathbb{R}^{3}$ that is non-zero only on dynamic objects if the ego-motion is correctly estimated.
III-A Method overview
@@
The method is summarized by pseudo-code in Algorithm 1. The proposed method first initializes two sets of clusters: hard and soft. Hard rigid clusters are non-overlapping small clusters that are assumed to cover only a single rigid object or its part. Each rigid cluster $H\in\mathcal{H}$ is a small compact cluster delivered through spatial-temporal segmentation on $P$, including the adjacent temporal point clouds. In contrast, the soft rigid clusters are small overlapping clusters that are expected to overflow into neighboring objects. Each point $\mathbf{p}\in P$ is associated with the one soft cluster $S\in\mathcal{S}$, which is defined as the set of its $k$ nearest neighbors $N_{k}(\mathbf{p},P)$ from the point cloud $P$. Given these two sets of clusters, we optimize the flow to minimize (i) hard rigidity loss on hard clusters, (ii) soft rigidity loss on soft clusters, and (iii) distance loss on all points. Both rigidity losses enforce rigid flow on points within the cluster. The main difference is that the soft rigidity loss allows for outlier rejection. All losses are detailed in the following paragraphs. Once the optimized flow is available, we merge rigid clusters in $P$, whose flow goes into the same rigid cluster in $Q$. Since the algorithm is iteratively called on all consecutive pairs of point clouds, the rigidity is propagated throughout the temporal domain.
@@
    Minimize
    $\mathcal{L}=\alpha~{}\sum_{\mathbf{f}\in F}\mathcal{L}_{\textrm{dist}}(\mathbf{f})+\beta\sum_{H\in\mathcal{H}}\sum_{\mathbf{f}\in H}\mathcal{L}_{\textrm{hard}}(\mathbf{f})+\gamma\sum_{S\in\mathcal{S}}\sum_{\mathbf{f}\in S}\mathcal{L}_{\textrm{soft}}(\mathbf{f}),$
    where $\alpha,\beta$ and $\gamma$ are hyper-parameters of the proposed method and $\mathbf{p}$, $\mathbf{q}$ and $\mathbf{f}$ are input points and corresponding flows. For simplicity, we omit per-point $\mathbf{p}$ and $\mathbf{q}$ notation.
@@
    Repeat from 3 until convergence or reaching a maximum number of iterations.
III-B Distance loss
@@
$\mathcal{L}_{\mathrm{dist}}=\sum_{i\in P}\|\mathbf{p}_{i}+\mathbf{f}_{i}-N_{1}(\mathbf{p}_{i}+\mathbf{f}_{i},Q)\|$ (1)
@@
III-C Hard rigidity loss